OpenStudy (anonymous):

Find the specific solution of the differential equation dy /dx = 4y/x^2 with condition y(−4) = 1. a: y =- 1 - 4/ x b: y =-e^(1/x) c: y=e^(-4/x) d: None of these

OpenStudy (anonymous):

@rational can u plz help? :)

OpenStudy (anonymous):

@nirmalnema

OpenStudy (anonymous):

@apoorvk

OpenStudy (nirmalnema):

dy /dx = 4y/x^2 dy/4y=dx/x^2 now u can easily int both the side.1!

OpenStudy (anonymous):

so what would i get as an answer? :)

OpenStudy (nirmalnema):

first u try ..then i will help u.

OpenStudy (anonymous):

i dont know how to integrate both sides :(

OpenStudy (nirmalnema):

simply int one by one LHS and RHS

OpenStudy (anonymous):

so would it be: 4y/x^2 & i have to integrate that?

OpenStudy (nirmalnema):

dy/4y=dx/x^2 int this.

OpenStudy (anonymous):

would it be: y(x) = c1 e^(-4/x)

OpenStudy (nirmalnema):

yup..!!

OpenStudy (anonymous):

so i think its C, y = e^(-4/x)

OpenStudy (nirmalnema):

yaa..!!

OpenStudy (anonymous):

thank u!!!!!!!! :)